OpenStudy (anonymous):

f(x)=square root of 11-x

OpenStudy (anonymous):

\[f(x)=\sqrt{11-x}\]what do you have to do?

OpenStudy (anonymous):

@fbi2015 Sorry! You have to determine the function.

OpenStudy (anonymous):

Idk what you meant by that, but...

OpenStudy (anonymous):

x-intercept\[0=\sqrt{11-x}~~~~take~~a~~squ are~~root~~of~~both~~sides.\]\[0=x-11~~~~~~x=11\]so x-intercept=11

OpenStudy (anonymous):

y-intercept\[y=\sqrt{11-(0)}\]\[y=\sqrt{11}~~~~this~~is~~the~~y~~intercept\]